What is the area and perimeter of 81ft by 39ft
Answer by JulietG(1812)

The problem gives you the following information:
L = 81
W = 39
To find the area, multiply the width by the length (A=L*W)
81 * 39 = 3159
Remember that perimeter is a fence. You'll need 2 pieces for the length and 2 for the width. (P=2L+2W)
(2*81)+(2*39)
162 + 78 =240